Solve the first order linear inhomogeneous differential equation using the Bernoulli method:

y,+(3y/x)=(1/x3)


1
Expert's answer
2022-01-06T06:32:26-0500
"y'+\dfrac{3}{x}y=\dfrac{1}{x^3}"

Let "v(x)=\dfrac{1}{x^2}"


"v'(x)=-\dfrac{2}{x^3}"

"-\dfrac{2}{x^3}+\dfrac{3}{x^3}=\dfrac{1}{x^3}"

The function "v(x)=\dfrac{1}{x^2}" is the solution of the given differential equation.

Let "y=u(x)v(x)=u(x)(\dfrac{1}{x^2})." Then


"y'=\dfrac{u'}{x^2}-\dfrac{2u}{x^3}"

Substitute


"\dfrac{u'}{x^2}-\dfrac{2u}{x^3}+\dfrac{3u}{x^3}=\dfrac{1}{x^3}"

"xu'+u=1"

"\dfrac{du}{u-1}=-\dfrac{dx}{x}"

Integrate


"\int \dfrac{du}{u-1}=-\int\dfrac{dx}{x}"

"u-1=\dfrac{C}{x}"

"y=(\dfrac{C}{x}+1)(\dfrac{1}{x^2})"

"y(x)=\dfrac{1}{x^2}+\dfrac{C}{x^3}"